### v1

Abstract
Computational notebooks have become the primary coding environment for data scientists. Despite their popularity, research on the code quality of these notebooks is still in its infancy, and the code shared in these notebooks is often of poor quality. Considering the importance of maintenance and reusability, it is crucial to pay attention to the understandability of the notebook code and identify the notebook metrics that play a significant role in their understandability. The level of code understandability is a qualitative variable closely associated with the user’s opinion about the code. Traditional approaches to measuring it either use limited questionnaires to review a few code pieces or rely on metadata such as likes and votes in software repositories. In our approach, we enhanced the measurement of the understandability level of Jupyter notebooks by leveraging user comments related to code understandability within a software repository. As a case study, we started with 542,051 Kaggle Jupyter notebooks, named DistilKaggle, which were introduced in our previous research. To identify user comments associated with code understandability, we utilized a fine-tuned DistilBERT transformer. We established a user opinion-based criterion for measuring code understandability by considering the number of code understandability-related comments, the upvotes on those comments, the total views of the notebook, and the total upvotes received by the notebook. We refer to this criterion as User Opinion Code Understandability (UOCU), which has proven much more effective than previous approaches. In addition, we were able to train machine learning models which could predict code understandability of notebooks solely based on their metrics. We collected a total of 34 metrics for 132,723 final notebooks as features in our dataset, while UOCU is considered as the label for them. Using the Random Forest classifier, our predictive model achieved 89% accuracy in predicting code understandability levels in computational notebooks.
Keywords:
Computational notebook Code understandability Kaggle Jupyter notebook metrics Machine learning
1 Introduction
Computational notebooks have become the primary coding environment for data scientists psallidas2019data . They offer several advantages over traditional software development environments, including easy documentation, code sharing, result analysis, visual and intuitive code development, and the ability to compile, execute, and modify code cells without re-executing the entire notebook. The advantages outlined above establish computational notebooks as a fundamental tool in the development of data analysis solutions. For instance, a case study at Microsoft highlighted that data scientists rely on computational notebooks for sharing and reusing their code across software teams epperson2022strategies . In another empirical study, cleaned computational notebooks were introduced as a best practice for a friction-less transition of data-driven projects from the exploration stage to production as a software engineering product quaranta2022assessing .
The quality and understandability of implementations in computational notebooks are crucial for various purposes, such as education to provide coding best practices for future data scientists wang2020better , notebook reusability epperson2022strategies , and maintainability wang2020assessing ; yang2021subtle ; dong2021splitting . Recent studies have focused on improving the quality of computational notebooks through enhanced documentation liu2019understanding ; rule2018exploration ; wang2022documentation ; yang2021subtle ; venkatesh2023enhancing , notebook structure wenskovitch2019albireo ; titov2022resplit ; jiang2022elevating , and managing notebook variants and revisions subramanian2019supporting ; jiang2022elevating . However, efforts to enhance code understandability in computational notebooks are still in the early stages venkatesh2021automated . This paper aims to address this problem.
Code Understandability (CU), also referred to as program understandability or code comprehensibility, is the active process by which developers acquire knowledge about a software system through exploring software artifacts, reading source code, and consulting documentation xia2017measuring . On average, developers spend approximately 60% of their time engaged in program comprehension during software development and maintenance xia2017measuring . To measure CU, various approaches have been proposed, often involving gathering feedback from individuals regarding a program or specific code segments. These studies employ different response variables, including time, accuracy, opinion, and visual metrics, as criteria for assessing CU. Some studies directly solicit opinions from a limited number of developers about specific code pieces sykes1983effect ; buse2009learning ; medeiros2018investigating ; scalabrino2019automatically . Other studies rely on metrics such as likes, stars, or votes on software repositories to gauge CU lu2018internal ; nasehi2012makes ; wang2022documentation ; liu2021haconvgnn . However, findings from small-scale studies may not be widely applicable or provide a comprehensive understanding of how developers perceive code. Additionally, user votes on software projects may be influenced by factors beyond CU, such as ease of use, algorithm accuracy, popularity, and marketing considerations.
In this paper, we have employed the opinion response variable oliveira2020evaluating to enhance the measurement of CU in a unique manner. Instead of conducting surveys, direct inquiries, or relying solely on user votes, we have developed a novel approach that involves extracting and analyzing comments regarding the understandability of notebooks in a repository. As a case study, we chose Kaggle [^1] https://www.kaggle.com, which is a notebook repository with a variety of interactive and collaborative functionalities. In this regard, we began by obtaining the metadata files of the Jupyter notebooks and then applied preprocessing to the user comments. Using a fine-tuned DistilBERT transformer, we identified user comments related to CU and provided a metric considering how much each comment attracted other users’ attention. We named this metric User Opinion Code Understandability (UOCU). Compared to a dataset labeled by 42 experts, UOCU has been shown to outperform current state-of-the-art methods for measuring the understandability of notebooks, as demonstrated in our evaluation. Incorporating the criterion of notebook upvotes in UOCU metric yields demonstrably superior outcomes in our evaluations, as evidenced by our research findings.
Another contribution of our study is the exploration of notebook metrics that have a more significant impact on CU. Previous studies have introduced a set of code metrics (such as lines of code, maximum line length, and cyclomatic complexity) that are related to CU for traditional programming languages like Java, Python, and C buse2009learning ; posnett2011simpler ; kasto2013measuring ; scalabrino2019automatically ; grotov2022large . However, some of these metrics (such as coupling, cohesion, and depth of inheritance tree) are not relevant to notebooks. Notebooks may have more metrics due to structural differences compared to code scripts addressed in recent studies grotov2022large ; wang2020assessing ; wang2020better ; titov2022resplit ; ono2021interactive ; wang2021restoring ; zhu2021restoring ; dilhara2021understanding ; titov2022resplit ; yang2021subtle ; wang2022documentation ; venkatesh2023enhancing . Additionally, studies that propose some notebook metrics as effective measures for assessing CU lack a systematic approach supported by significant big data.
To address this gap, we conducted a comprehensive study of Jupyter notebook metrics to identify 34 metrics related to CU. We employed a machine learning approach to investigate the correlation between CU and Jupyter notebook metrics. For this purpose, we leveraged the DistilKaggle dataset, which comprised 542,051 notebooks shared on the Kaggle platform and was originally introduced in mostafavi2024DistilKaggle . We preprocessed KGTorrent data and distilled a dataset, extracting 34 metrics from each notebook. Finally, we merged this dataset with the output pertaining to UOCU, allowing us to analyze the relationship between the metrics and CU in a comprehensive manner.
After assembling a dataset containing notebook metrics and their associated CU scores, we developed a machine learning model using four different algorithms, including CatBoost, XGBoost, Decision Tree, and Random Forest. The best result we achieved for predicting the understandability of notebook codes was with the Random Forest algorithm, which achieved an accuracy of 89% and an F1-score of 88%, improving previous criteria. With this acceptable level of accuracy and F1-score, we identified the most influential notebook metrics for CU. These metrics included developer expertise level, number of blank lines of code, nubmer of visualization data types, number of executed cells, number of H1 tags in markdown headlines and number of lines of markdown cells.
In summary, our contributions are as follows:
  • creating a dataset for CU through fine-grained elicitation of understandability-related metrics
  • introducing a new criterion for CU of computational notebooks
  • generating a dataset and creating a model to predict the understandability levels within Jupyter notebooks
  • identifying the most impactful notebook and code metrics for improving the CU
This paper is organized as follows: Section 2 presents a motivating example. Section 3 provides a systematic review of notebook metrics. Section 4 identifies the problems and research questions in detail. Our proposed approach is described in Section 5, while Section 6 presents our evaluations and results. Section 7 discusses related work. The limitations and validity risks associated with our findings are explained in Section 8. Finally, Section 9 concludes the paper and provides future research directions.
Our dataset and experimental source code are available online [^2] https://github.com/ISE-Research/NotebookCU. Moreover, our service, which accepts a notebook and returns its code metrics and predicts its CU level is accessible [^3] http://46.249.101.150/docs.
2 Motivating Example
Some studies leverage the quantity of user votes received by a code or notebook within a software repository as an indicator of its understandability level nasehi2012makes ; lu2018internal ; wang2022documentation ; liu2021haconvgnn . However, user votes can be influenced by various factors such as output accuracy, popularity, and ease of reuse, among others. We have selected 10 notebook samples from the Kaggle repository along with users’ upvote data up to August 10, 2023 (latest upvote in Table 1). Three programming experts with at least three years of experience working with Jupyter notebooks were asked to evaluate each example and place them into either a normal or good CU group. Table 1 illustrates the average outcomes of the experts’ evaluations, representing the prevailing expert assessment of CU across the notebook samples. As can be seen, in these notebooks, the experts’ opinions contradicted the results of the user votes. Upon an initial review of the comments related to these notebooks, it is evident that the comments for the first four notebooks have little relevance to CU. Conversely, the comments regarding CU for the last four notebooks (e.g.,“nicely organized file and good EDA”, “well-explained notebook”, “very well presented and nice kernel”) are in a significantly better state, based on the number of visitors and upvotes received. Therefore, compared to the opinions of experts, these comments are considered to reflect a higher level of understandability satisfaction.
Table 1: Examples of Notebooks Upvotes and Experts’ Scores
Notebooks* Latest Upvote Expert Score
A 66 normal
B 72 normal
C 96 normal
D 113 normal
E 5 good
F 21 good
G 21 good
H 40 good
* Links of notebooks:
    A:https://www.kaggle.com/code/panamby/loan-default-prediction
    B:https://www.kaggle.com/code/christofhenkel/weighted-kappa-loss-for-keras-tensorflow
    C:https://www.kaggle.com/code/bicotsp/pmtest1
    D:https://www.kaggle.com/code/raddar/paris-madness
    E:https://www.kaggle.com/code/naren204/speed-at-which-pandemic-covid-19-spreading
    F:https://www.kaggle.com/code/avnika22/lgbm-predicting-house-prices
    G:https://www.kaggle.com/code/nikhileshkos/play-store-analysis
    H:https://www.kaggle.com/code/landfallmotto/60k-stack-overflow-questions-keras-lstm-and-cnn
Table 2: Examples of Notebooks and their Metrics
ID
Lines of Code
Lines of Comment
Num. of Code Cells
Num. of MD* Cells
Num. of Executed Cells
Num.of Visual Data
Num.of Headlines
Performance Tier
level of CU
A 239 69 34 11 16 1 1 1 normal
B 38 0 2 8 0 0 0 4 normal
C 166 10 27 1 11 0 0 0 normal
D 405 6 76 21 5 2 5 4 normal
E 214 13 44 28 16 10 12 0 good
F 191 2 48 46 24 8 8 2 good
G 99 15 32 19 46 15 17 2 good
H 140 12 22 38 18 5 15 2 good
* MD: markdown
Furthermore, our objective is to identify notebook metrics that have a significant impact on CU. To achieve this, we have calculated a subset of code metrics expected to serve as an effective indicator of CU. Table 2 displays the calculated metrics for these notebooks. The process of choosing and calculating them, as well as how their impact on CU was identified, is described in Section 3.
Based on the expert opinions in Table 1 and the metrics provided in Table 2, we can hastily conclude that there is a direct relationship between some notebook metrics and CU. Specifically, sufficient documentation using markdowns and headlines, preserving the output of each code cell, and presenting results using visualization tools all contribute to a better understanding of notebook code. In this paper we aim to explore the applicability of these findings to other cases and examine additional metrics in order to assess their generalizability.
3 Metrics of Notebooks
Metrics play a vital role in evaluating the effectiveness, comprehensibility, and quality of computational notebooks. Metrics provide a systematic framework for measuring and analyzing various aspects of notebooks, enabling a thorough assessment. In this section, we delve into two distinct groups of metrics: Notebook-Based Metrics and Script-Based Metrics. Notebook-Based Metrics focus on the properties and structure of the notebook itself, including organization, documentation, and visual representation. On the other hand, Script-Based Metrics concentrate on the underlying code, examining factors such as complexity, modularity, and adherence to coding standards. Through the exploration and analysis of these metrics, our aim is to gain deeper insights into the strengths, weaknesses, and overall effectiveness of computational notebooks.
3.1 Notebook-Based Metrics
Our first set of metrics focuses on assessing the quality of Jupyter notebooks, a topic that has received considerable attention in recent years. Numerous articles have explored the evaluation of code quality in this environment, introducing a variety of metrics grotov2022large ; wang2020assessing ; wang2020better ; titov2022resplit ; ono2021interactive ; wang2021restoring ; zhu2021restoring ; dilhara2021understanding ; yang2021subtle ; wang2022documentation ; venkatesh2023enhancing ; liu2021haconvgnn . We conducted an extensive review of literature from top-tier journals and conferences in software engineering and data science domains, employing snowballing techniques to augment our search and uncover additional relevant contributions. In selecting metrics related to CU (CU), our approach was guided by the comprehensive definition provided by Xia et al. (2017), emphasizing the active process through which developers acquire knowledge about software systems, particularly within the Jupyter notebook environment. This definition underscores activities such as exploring software artifacts, reading source code, and consulting documentation. Leveraging insights from domain experts with extensive expertise in Python development, averaging 5 years of experience in software engineering or data science, we refined an initial list of 23 metrics down to 17, prioritizing those most pertinent to CU in Jupyter notebooks. These metrics are detailed in Table 3. Subsequently, we will elaborate on metrics whose names may not clearly indicate their conceptual underpinnings.
[b] Feature Abbreviation Reference Script-Based Metrics Basic Metrics Lines of Code LOC  scalabrino2019automatically ; buse2009learning ; kasto2013measuring ; wang2022documentation Number of Blank Lines of Code BLC  scalabrino2019automatically ; buse2009learning ; kasto2013measuring Lines of Comments LOCom  scalabrino2019automatically ; buse2009learning ; kasto2013measuring Number of Comment Words CW  kasto2013measuring Number of Statements S  kasto2013measuring ; scalabrino2019automatically Number of Parameters P  kasto2013measuring ; scalabrino2019automatically Number of User-Defined Functions UDF  grotov2022large Complexity Metrics Cyclomatic Complexity CyC  scalabrino2019automatically ; buse2009learning ; kasto2013measuring ; grotov2022large Nested Block Depth NBD  kasto2013measuring ; scalabrino2019automatically Kind of Line of Code Identifier Density KLCID  klemola2003cognitive Halstead Metrics Number of Operands OPRND  posnett2011simpler ; kasto2013measuring ; scalabrino2019automatically Number of Operators OPRAT  posnett2011simpler ; kasto2013measuring ; buse2009learning Number of Unique Operands UOPRND  posnett2011simpler ; kasto2013measuring Number of Unique Operators UOPRAT  posnett2011simpler ; kasto2013measuring Readability Metrics Average Line Length of Code ALLC  buse2009learning Number of Identifiers ID  buse2009learning Average Number of Identifiers AID  buse2009learning Notebook-Based Metrics Number of Code Cells CC  pimentel2019large ; wang2022documentation Mean Number of Lines in Code Cells MeanLCC  kasto2013measuring Number of Imports I  pimentel2019large ; kasto2013measuring Mean Number of Words in Markdown Cells MeanWMC  wang2022documentation Number of H1 tags in the Markdown H1  pimentel2019large ; venkatesh2021automated ; venkatesh2023enhancing Number of H2 tags in the Markdown H2  pimentel2019large ; wang2020assessing ; venkatesh2023enhancing Number of H3 tags in the Markdown H3  pimentel2019large ; wang2020assessing ; venkatesh2023enhancing Number of Markdown Cells MC  wang2022documentation ; pimentel2019large ; liu2021haconvgnn Mean Number of Lines in Markdown Cells MeanLMC  grotov2022large Number of Markdown Words MW  kasto2013measuring ; yang2021subtle Number of Lines in Markdown Cells LMC  scalabrino2019automatically ; wang2022documentation Distribution of Markdown Cells DMC  pimentel2019large ; wang2022documentation Distribution of Imports DI  pimentel2019large ; wang2022documentation Performance Tier PT  alomar2021behind ; scalabrino2019automatically External API Popularity EAP  scalabrino2019automatically ; zhu2021restoring ; dilhara2021understanding Number of Visualization Data Type VDT  ono2021interactive ; agrawal2022understanding Number of Executed Cells EC  pimentel2019large
Table 3: Features Extracted in this Study
  • Number of Visualization Data Type: Considering that it is possible to display the output of each cell in notebooks by executing it, this metric shows the number of visual outputs, including all kinds of images and graphs.
  • Number of Executed Cells: This attribute specifies the number of cells that have been executed in a notebook and have an execution order number.
  • H1, H2 and H3 headlines: These metrics show the number of headlines that are written inside the markdowns and are usually used to express titles and subtitles.
  • Distribution of Markdown Cells: For calculating the distribution of markdowns throughout a notebook, we suggest the inverse coefficient of variation. The coefficient of variation (CV), also known as relative standard deviation (RSD), is a statistical measure commonly used for comparing diversity in workgroups and is defined as the ratio of the standard deviation to the mean bedeian2000use .
    $CV=\sigma/\mu$ (1)
    where:
    $\sigma=\sqrt{\sum(x_{i}-\mu)^{2}/n}$ (2)
    where:
    - n: the size of population
    - $x_{i}$: each value from the population
    To calculate the distribution of markdown cells, we considered the number of words of each markdown as the size of it. After applying the inverse of Equation (1) to the notebooks of the final dataset, the highest CV (the most unbalanced distribution) assigned to notebook P [^4] P:https://www.kaggle.com/code/xiwuhan/jmtc-2nd-place-solution and the lowest CV (the most balanced distribution) was for notebook Q [^5] Q:https://www.kaggle.com/code/anokas/two-sigma-time-travel-eda.
  • Distribution of Imports: To calculate the amount of distribution, we used Equation (1), too. Here, we consider the size of each code cell as the number of imports in it. In this way, the highest CV was assigned to notebook R [^6] R:https://www.kaggle.com/code/moazmagdy/titanic-survival-prediction, in which, despite the high number of code cells, all import commands were accumulated in one cell and the lowest CV was assigned to S [^7] S:https://www.kaggle.com/code/arturkiulian/coronawhy-org-team-980-people-join-slack as it only contained one import command in each code cell.
  • Performance Tier: This feature is present in the metadata of the Kaggle repository, which, based on specific rules, determines the level of expertise of notebook developers with one of the levels 0 to 4888https://www.kaggle.com/progression. Details of how Kaggle calculates the performance tier are given in Appendix B.
  • External API Popularity: This metric assigns a number to each notebook that shows to what extent the APIs and libraries used in that notebook are popular zhu2021restoring . The higher its value, the more frequently other developers utilize the APIs featured in the notebook, which results in better CU based on previous studies scalabrino2019automatically . In order to measure this criterion in notebooks: First, we counted the number of times each API was imported in our dataset and assigned a popularity score to each API based on the frequency at which they were used in the whole dataset. Then, for each notebook, we summed up the popularity score of APIs used in that notebook to obtain the External API popularity score for that notebook.
3.2 Script-Based Metrics
Given that each code cell in Jupyter notebooks is a regular Python script, many of the metrics for code scripts proposed by prior studies buse2009learning ; dorn2012general ; kasto2013measuring ; scalabrino2019automatically ; mathias1999role ; parker2003measuring are also applicable to code cells in Jupyter notebooks. Considering that a small percentage of notebooks use the concept of class and object orientation [^9] Based on the statistics of our notebooks dataset, only 3.5% of notebooks defined classes to implement objects. and that notebooks typically have a weaker structure, some metrics were ignored.
Script-based metrics include seven basic metrics, three complexity metrics, four Halstead metrics and three readability metrics as introduced in Table 3. Some metrics that require further explanation will be mentioned here:
  • Basic metrics: Basic metrics that are usually considered in most related articles are: the number of lines of code, number of blank lines of code, number of comment lines of code, number of comment words, number of statements, number of parameters and number of user-defined function buse2009learning ; dorn2012general ; kasto2013measuring ; scalabrino2019automatically .
  • Halstead metrics: Halstead metrics were employed by Parker and Becker parker2003measuring for the first time, to measure and compare the effectiveness of students’ solutions of two different code writing assessments based on the premise that the metrics can be seen as a measure of work done. Kasto kasto2013measuring introduced 4 of Halstead’s metrics as imperative metrics, which are: the number of operands, number of operators, number of unique operands and number of unique operators.
  • Cyclomatic Complexity: Cyclomatic complexity assigns a number for code complexity based on the code graph. It was first used by Mathias mathias1999role in order to examine the underlying nature of code designed to study the process of program comprehension.
  • KLCID: Kind of Line of Code Identifier Density (KLCID) klemola2003cognitive is one of the complexity metrics that analyzes the cognitive complexity of program comprehension tasks. The KLCID is calculated by counting the lines of conceptually unique code and calculating the density of identifiers. The KLCID computation involves transforming code into a format where lines with similar structures are identified. Unique lines with identifiers are counted, reflecting distinct patterns. KLCID is obtained by dividing the number of identifiers (such as variables, function names, or constants) by total lines, aiming to assess cognitive complexity. It recognizes that lines with more identifiers are harder to understand, emphasizing comprehension challenges posed by dense identifier usage in code segments.
4 Problem Definition
Given the challenges posed by the limited understandability of Jupyter notebooks and the deficiencies in existing methods for assessing CU within this context, this paper presents a pioneering methodology for quantifying CU in Jupyter notebooks. Furthermore, it identifies the primary notebook metrics that significantly impact CU. To clarify the problem, in this section, we raise three research questions. Subsequently, we present our proposed approach (Section 5) and evaluate its effectiveness (Section 6) in addressing these questions.
RQ1: How accurately is it possible to provide a criterion for measuring code understandability of computational notebooks based on user comments?
In numerous studies sykes1983effect ; buse2009learning ; medeiros2018investigating ; scalabrino2019automatically , expert opinion has been employed as a criterion to measure CU. Questionnaires have been utilized as tools to assess the comprehension of experts based on their opinions regarding specific code snippets. In our study, we have also incorporated the criterion of expert opinion. However, our approach is unique as we analyze comments and opinions shared by individuals in notebook repositories. Nevertheless, two vital issues need to be addressed regarding these comments. Firstly, we must determine which comments are directly relevant to CU. Secondly, we need to assess the extent to which these comments have received attention from other experts. Taking these aspects into account, it is essential to define a criterion that effectively measures the ability to understand notebooks. Evaluating whether this criterion can independently assess CU or complement the previously established criteria is important. By comparing the results obtained from this criterion with a baseline, we can determine whether the new approach represents an improvement over existing methods, establishing its effectiveness and reliability.
RQ2: Is it feasible to develop a model capable of achieving satisfactory performance in assessing the code understandability of notebooks?
Previous studies, such as scalabrino2019automatically ; buse2009learning have focused on developing methods for automatically measuring CU. However, these studies either did not provide a specific model or their model was limited to Java-based programs. In contrast, the primary objective of this study is to develop a model specifically for notebooks that can accurately estimate CU. In order to develop this model, two factors must be taken into consideration: the code metrics of the relevant notebooks for CU, and a suitable criterion for assessing CU based on the findings from RQ1. Previous studies have utilized upvotes as a metric to gauge the CU of large-scale notebooks. Therefore, a model trained using this criterion can serve as a valuable benchmark for comparison with our own model. If our method demonstrably outperforms previous methods on evaluation parameters, it can serve as an initial step toward addressing RQ3 in a positive manner.
RQ3: Which notebook metrics can affect CU, and to what extent can each metric influence code understandability?
After identifying the metrics related to CU and learning how to extract them (as described in Section 3), answering this research question requires addressing the following questions: How can we measure the impact of these metrics on notebook CU? Which metrics are the most effective in improving its understandability? Which group of metrics prove to be the most impactful in improving CU of notebooks? By addressing these research questions, we aim to pinpoint script-based and notebook-based metrics that could either enhance or hinder CU. By pinpointing the most influential metrics or groups of metrics, we can glean valuable insights on how to craft more comprehensible notebooks, thus fostering improved programming practices.
5 The Proposed Approach
The proposed approach consists of two key components. The first component of the proposed approach involves developing a criterion to assess the understandability of notebook code in Sections 5.1-5.2. This criterion was effective in addressing RQ1.
Figure 1: Proposed Approach Overview to Predict Two Class of Notebooks: Normal Code Understandability (NCU) and Good Code Understandability (GCU).
The second component involves preparing a comprehensive dataset of notebooks and generating the metrics shown in Table 3. Different machine learning algorithms are then applied to find the best algorithm to determine the most effective metrics on CU. This analysis takes place in Sections 5.3-5.4.
In fact, in Section 5.4, machine learning algorithms are used to find answers to RQ2 and RQ3. If we consider the learning operation as y=f(X), the output of Section 5.3 is used as X, and the output of Section 5.2 is used as y.
The machine learning algorithms in Section 5.4 utilize the findings from Section 5.3 as the input (X) and the results from Section 5.2 as the desired output (y). These algorithms aim is to find a function f (y=f(x)) that maps the input data X to the desired output y. This function can then be used to predict new, unseen data. Utilizing this learning operation, the algorithms aim to discover patterns or relationships between the notebooks’ code metrics and their CU, which can help provide answers to RQ2 and RQ3.
Figure 1 shows our approach’s steps to answer the research questions, which we will discuss in detail about each step.
5.1 Identifying Comments Related to CU
The comments corresponding to each notebook in the DistilKaggle dataset quaranta2021kgtorrent are available using Kaggle metadata [^10] https://www.kaggle.com/datasets/kaggle/meta-kaggle. The large number of comments available in our dataset gave us the opportunity to conduct an exploratory study and analyze these comments with the aim of establishing a metric for Jupyter notebooks’ CU.
During the data cleaning stage, any non-English comments were removed. We encountered a decline in performance during the training of our deep learning models, primarily attributed to the presence of lengthy comments within our dataset. Consequently, by filtering out comments longer than the 99th percentile of the dataset’s length distribution, we significantly reduced our training time while retaining the majority of our data. These comments had between 1,000 and 17,000 characters, which is less than 1% of the total data records.
Furthermore, comments posted by the authors of the notebooks themselves were excluded from the analysis. Similarly, comments that were shared in response to others’ comments were also excluded. These types of responses often consist of simple expressions of gratitude or acknowledgments, which do not provide meaningful insights into the understandability of the notebook.
In the following, we outline our methodology for classifying user comments relating to Jupyter notebooks on Kaggle into three categories: “positive in terms of understandability,” “negative in terms of understandability,” or “no relation to understandability,” utilizing a team of annotators. Initially, we crafted a guideline outlining definitions for each category supplemented with illustrative examples. This guideline served to aid annotators in comprehending the process more effectively, thereby reducing discrepancies. We had a total of 659,894 comments. Based on calculations similar to those outlined by Zhu et al. zhu2021restoring , we determined that a sample size of 1,039 comments was required to attain a confidence level of 99% with a margin of error of ±4% sample2024 . Consequently, about 1,000 comments were extracted to train our model.
Four authors of this paper, having expertise in CU concepts, were selected as annotators for the labeling procedure. Each comment underwent assessment by two annotators who conducted an in-depth manual analysis of the content and assigned appropriate labels. An overall consensus was attained on 92% of the labels, resulting in a Cohen’s kappa coefficient of 80%, signifying reliable outcomes cohen1960coefficient . All contentious comments were meticulously reviewed, and a collaborative session was conducted with annotators to finalize the labeling decisions. Annotators engaged in productive discussions to harmonize and unanimously agree on the definitive labels.
After analyzing the labels, it was noted that there were six negative comments, which accounted for less than 1% of the total comments. This pattern may have been influenced by the competitive nature among members participating in various data science competitions, ensuring that notebooks meet specific clarity and quality standards.
The distribution of these three groups of comments is shown in Figure 2. For a better understanding of how to label comments, several examples of comments and their labels are given in Appendix A.
Figure 2: Distribution of User Comments
Considering this observation, we removed negative comments and consolidated the labels for each comment into two categories: “related to code understandability” (which are positive comments) and “unrelated to code understandability”. This distribution of comments motivates our paper’s exploration of an intelligent approach for labeling notebooks as “normal” or “good” in terms of code understandability, as detailed in the Section 5.2.
Therefore, out of the total 994 results, 467 were categorized as “related to code understandability” and 527 as “unrelated to code understandability”. To pinpoint comments specifically relevant to CU within the entire dataset, we fine-tuned the DistilBERT  [^11] DistilBERT is a smaller general-purpose language representation model and a distilled version of BERT, which is smaller, faster, computationally cheaper, and lighter. It can be used for text classification, named entity recognition, question answering, text generation, machine translation, and recommender systems. model on this data. DistilBERT was chosen due to its ability to effectively capture the contextual semantics of full sentences, which is crucial for understanding user comments. Additionally, as a distilled version of BERT, DistilBERT offers a good balance between performance and efficiency, making it suitable for our task. It is pre-trained on a large corpus of text, allowing it to understand the context of a sentence, and can be fine-tuned on a specific task, like classifying comments in this case. We had two main objectives in training this model. Firstly, we wanted to filter out comments that were unrelated to CU in order to evaluate the notebooks based only on their relevant comments. Secondly, we aimed to identify and include all comments related to CU when assigning an understandability score to the notebooks. The model achieved an F1-score of 90% and  accuracy of 92%, identifying 410,484 out of 659,894 comments in the whole dataset that were related to CU. Finally, a score of one was assigned to comments related to CU, while a score of zero was assigned to other comments.
5.2 Calculating Notebook Score
To calculate the CU score of notebooks, we did two main tasks, which are presented in two sub-sections as follows.
5.2.1 Calculating CU Scores
As mentioned, the UOCU metric is utilized to measuring CU based on users’ opinions. To calculate UOCU for each notebook, we use Equation (3) which involves a function f that takes some metadata of a notebook and the BERTScore calculated in Section 5.1. This function then returns a new score called the “comment-based score”.
$UOCU=[\sum_{i=1}^{n}(BERTScore_{i}*(CommentVote_{i}+1))+\alpha]/\sqrt{TotalView}$ (3)
where:
- n is the total comments of a notebook,
- BERTScore is the binary score calculated for a notebook with DistilBERT,
- CommentVote is the number of upvotes for each comment,
- TotalView is all views of a notebook.
Currently, DistilBERT’s model generates a score for each comment that is either 1 (related to CU) or 0 (unrelated to CU). The details of Equation (3) are explained below.
  • Apart from the relevance of a comment to the CU, the votes given to that comment show users’ agreement. To augment this impact, we multiply it by “(CommentVote + 1)”. Each comment starts with a value of one (+1), and with each additional upvote, this value increments by one.
  • We calculated the total scores ($\Sigma$) of comments for each notebook, which yielded the initial notebook score.
  • We divided the final notebook scores by the square root of the number of times the notebook was viewed on Kaggle. This was done to mitigate the impact of views on notebooks. This adjustment ensures that notebooks with higher view counts (and consequently more comments) do not receive disproportionately large scores, while notebooks with fewer view counts (and potentially fewer comments) do not receive unfairly lower scores.
  • We have introduced a “+$\alpha$ numerator” to differentiate notebooks with varying view counts that may not have CU-related comments. As illustrated in Equation (3), excluding the “+$\alpha$” would result in a numerator of zero for notebooks with zero comments, rendering them indistinguishable. With the inclusion of “+$\alpha$” in the numerator, we generate a small value that is then divided by the square root of the total views, creating diverse numbers. Through iterative experimentation, we have consistently observed that a value of 5 yields superior results in accomplishing this objective.
5.2.2 Converting CU Scores to Binary Class Labels
Up to this point, we have effectively formulated the UOCU benchmark leveraging DistilBERT. Should this criterion prove to be more fitting for comprehending notebook code compared to previous criteria(which is proved in the Section 6), it can be adopted as the label for the dataset, enabling the training of a model to assess the understandability of notebooks.
In this section, our primary emphasis is on converting the continuous CU score into discrete binary labels. Although we assess machine learning performance using four-level and three-level class labels, the main focus remains on binary classification.
In the following two paragraphs, the reason behind the converting the CU score to discrete class labels is elaborated. Linear regression is perhaps the most commonly used tool in statistics and as such is familiar to a diverse range of students and researchers. However, an even wider segment of the educated public do not understand regression or least squares and thus have to take many statistical results on faith, for example accepting results that are labeled as statistically significant without being able to interpret their numerical values. As a result, we would like to approximate the regression using a simple comparison. For this purpose, we can simplify the interpretation of regression by discretizing. In common practice, variables are discretized into two categories (i.e., the predictor x falling above or below some threshold) Andrew2009Splitting . However, as shown in this study Andrew2009Splitting , we can do better by discretizing the objective variable into three values and throwing away the middle category. The study shows the efficiency of this method. It makes the general recommendation that the high and low categories each be set to contain 1/4 to 1/3 of the data, which results in comparisons with approximately 80%–90% of the efficiency of linear regression.
After finding the best scoring strategy and calculating the score for each notebook, we followed the same rule and discretized our score. We ranked notebooks according to their scores and divided them into four equally sized bins. Subsequently, we left out two bins in the middle and retained the first and the last quarters. The first quarter comprised notebooks with the lowest scores, labeled as Normal CU (NCU), and the last quarter encompassed notebooks with the highest scores, labeled as Good CU (GCU). Each of these selected bins contained 1/4 of the data.
5.3 Dataset
Various platforms offer large datasets of notebooks, providing a wealth of information about the notebooks themselves and their creators. Kaggle, a subsidiary of Google, is an online community specifically designed for data scientists and machine learning practitioners. Kaggle allows users to discover and publish datasets, explore and construct models within a web-based data science environment, collaborate with other professionals in the field, and participate in competitions to solve data science challenges. Similar to social media platforms, Kaggle incorporates features such as following a user and liking shared notebooks or datasets. The abundance of comments found under shared notebooks, which aid in CU, makes it valuable for employing opinion mining and extracting knowledge. Users have the ability to vote for a notebook or comment and endorse its content. Considering these attributes, we believe that any repository encompassing these features would be a suitable choice for implementing the approach outlined in this paper.
We initiated our study by utilizing the DistilKaggle dataset that was introduced in our prior research mostafavi2024DistilKaggle . This dataset offers a detailed insight into the content structure of 542,051 Jupyter notebooks, facilitating in-depth analysis of both code and markdown cells. Additionally, it included a notebook code metrics dataset that concentrated on identified code metrics of the notebooks. The methodology for constructing this tiny dataset is elaborated in Section 3.
5.4 Utilizing Machine Learning
We now have the features dataset and UOCU (as labels) ready to run a machine learning operation. The goal of this part is to discover if there is a relationship between the metrics of notebooks and their CU. It is also possible to provide a model that can estimate the level of understandability of a notebook. As the dataset features are all numbers and are not that complex, we employed four machine learning models to discover the relationship. We employed Random Forest, CatBoost, Decision Tree, and XGBoost models for the classification of notebooks into binary classes. CatBoost is a gradient boosting algorithm that handles categorical features efficiently, while XGBoost is known for its scalability, speed, and ability to handle missing data. Decision trees are interpretable and can capture complex relationships within the data. Finally, Random Forest combines multiple decision trees to improve accuracy and reduce overfitting, making it particularly useful when dealing with high-dimensional data and noisy features. By experimenting with these diverse algorithms, we aimed to find the best-performing model for our specific task of predicting notebook CU. Additionally, their simplicity allowed us to concentrate on the analysis of the features rather than intricate model architectures or complex hyperparameter tuning. The results of the evaluation metrics and the accuracy of these algorithms are presented in Section 6.2.3. The Random Forest algorithm ultimately stood out with an accuracy of 89% and F1-score of 88%. We conducted a grid search on a subset of the data to explore different hyperparameter configurations. We found that the default settings for these algorithms yielded competitive results while significantly reducing the computational overhead.
6 Evaluations
In this section, we will describe the experimental design that is needed to answer research questions. We will also evaluate our approach for generating notebook scores based on notebook comments on Kaggle. After ensuring the quality of notebook scores, we train our models based on these scores and evaluate their performance in different settings. The quantitative and qualitative results of these evaluations and their comparison with the baseline will be described in each section.
6.1 Experiment Design
To answer the questions of this research, we first set up the experiments. We trained our model on Google Colab, a cloud-based platform with a 12.7GB RAM, a 78GB Disk, a T4 GPU, and a Python 3 Engine. The important issues in the design of our experiments are described below.
6.1.1 Sampling
To streamline and optimize the study, a filter was applied to eliminate notebooks. After extensive experimentation, we observed that UOCU typically does not yield favorable scores for notebooks with fewer than 500 views. This phenomenon likely arises from the potential for an increase in comments as views accumulate. Therefore, we removed them from the 542,051 notebooks in the DistilKaggle dataset, resulting in a reduced count of 132,723 notebooks. Based on calculations similar to those outlined by Zhu et al. zhu2021restoring , we determined that a sample size of 1,050 notebooks is required to attain a confidence level of 99% with a margin of error of ±4% sample2024 . Consequently, these 1,050 notebooks were extracted and excluded from the training set and preserved to establish ground truth.
6.1.2 Human Labeling
We designed an experiment to assess the performance of the CU scores introduced in this paper. We conducted a human labeling task in which each annotator categorized notebooks into “good” and “normal” CU categories, following the definition provided by Xia et al. in Section 1.
Experts were assigned to annotate the sample dataset outlined in Section 6.1.1. Each notebook was evaluated by three experts and classified as either ‘NCU’ (normal code understandability) or ‘GCU’ (good code understandability).
To establish a ground truth dataset, we followed the labeling process methodology employed by Guzman et al. guzman2015ensemble . Initially, we developed a set of instructions containing definitions for each class and several examples showcasing various characteristics such as line count, comments, markdown, different coding styles, display outputs, etc. This guideline assisted annotators in understanding the classification process and helped minimize disagreements. Subsequently, a random selection of 1,050 notebooks detailed in Section 6.1.1 was made.
We invited 55 experts to participate, and eventually received responses from 42 of them. We used convenience sampling to select these experts. They were selected for inclusion in the sample because they were the easiest for us to access. Through the information gathered from these participants, we determined an average experience level of 5 years working with Python in scientific or industrial contexts. The notebook samples were systematically selected by the authors and randomly assigned to experts, each assessed by three annotators to ensure comprehensive evaluation and reliability of results. To ease the review process and conceal metadata like user comments, upvotes, and views, the notebooks were placed on nbviewer.com. The annotators conducted a manual analysis of the notebooks and assigned the appropriate labels. Subsequently, we addressed any disagreements in the reviews and conducted a meeting to reach final decisions, particularly for annotators with unlabeled notebooks.
In order to ensure the reliability and validity of our questionnaire, we employed kappa coefficients as a measure of inter-rater agreement. Given the involvement of three raters in our task, Cohen’s kappa was not applicable due to its restriction to two raters. Therefore, we utilized Fleiss’ kappa, which accommodates more than two raters. This approach allowed us to assess the consistency among multiple raters and provide robust assurance of the questionnaire’s quality. More details about Fleiss’ kappa are described in Section 6.1.4.
This comprehensive evaluation yielded a Fleiss’ kappa coefficient of 82%, providing almost perfect agreement of the questionnaire’s quality and reliability landis1977application . The labeled notebooks comprised 566 (54%) labeled as ‘NCU’ and 484 (46%) labeled as ‘GCU’.
6.1.3 Baselines
For our dataset, we needed to find models and criteria suitable for the specific task of CU classification of Jupyter notebooks. To accomplish this, we selected three different methods:
  1. 1.
    The upvote approach: We considered the number of upvotes on a notebook as a potential proxy measure for CU. As detailed in Section 7, certain studies that evaluate CU via user opinions have employed criteria such as the number of upvotes, likes, and stars to gauge this aspect within software repositories. Hence, we adopt this metric as a scoring mechanism for measuring CU within our dataset.
  2. 2.
    The CodeBERT codebert model fine-tuned on upvotes: In this method, we fine-tuned a transformer model on our dataset and used that as the baseline for our task.
  3. 3.
    The PLBART ahmad2021unified model fine-tuned on upvotes: Also in this approach, we fine-tuned another transformer model and used it as another baseline for our task.
CodeBERT has been pre-trained in both programming language (PL) and natural language (NL). It learns representations that are useful for various NL-PL applications such as natural language code search and code documentation generation. PLBART (Pretrained LAnguage- and Vision-based BART) is a state-of-the-art deep model that extends BART to incorporate both language and vision inputs, enabling it to generate high-quality text in multiple languages while understanding visual information. We employed the Hugging Face Transformers library for our work huggingface2019transformers . We did not have access to labeled data for fine-tuning these transformers. We only had 1,050 notebooks as our ground truth, which were used as our final test dataset to validate the performance of our methods. To address the scarcity of training data, we employed a straightforward approach of ranking notebooks based on their upvotes, considering it as a rudimentary metric for assessing CU. This approach aligns with previous studies, which often consider highly-voted notebooks as a proxy for well-documented notebooks Wang2021WhatMakes ; mondal2023cell2doc , or utilize the number of stars and forks as proxies for the quality of data science GitHub projects morakot2023mining ).
Therefore, we sorted the notebooks according to their upvotes. Subsequently, we curated our dataset by selecting an equal number of notebooks from both ends of the ranking spectrum – notebooks with the lowest scores (indicating minimal understandability score) and those with the highest scores (indicating significant understandability score). The specific number of notebooks chosen for inclusion in the dataset was determined experimentally.
Our analysis, as depicted in Figure 3, revealed a trend wherein increasing the sample size for fine-tuning baseline models led to improved performance, up to a threshold of 4,000 samples. Beyond this threshold, we observed no significant improvements in model training performance. Based on these findings, we concluded that a total sample size of 4,000 notebooks would be optimal for our study, comprising 2,000 notebooks with the lowest upvote scores, referred to as ’NCU,’ and an additional 2,000 notebooks with the highest scores, referred to as ’GCU.’
Figure 3: The Impact of Changing Sample Size on Fine-Tuning Baseline Models
We partitioned our dataset of 4,000 samples into training, validation, and test sets using a split ratio of 70-20-10. To tokenize the input text, we employed both the CodeBERT and PLBart base tokenizers, both derived from the RoBERTa tokenizer architecture. Our tokenization process involved separate tokenization of markdown and code segments, followed by truncation of each part. In our investigations with CodeBERT, we identified 350 characters as the optimal token limit for markdown, allocating the remaining space to the code segment. Similarly, for PLBart, we assigned 512 characters to markdown and allocated the remaining space to the code segment.
Regarding training parameters, we employed accuracy metric to assess model performance. We configured the “evaluation_strategy” as “epochs” and conducted training over 10 epochs. We restricted the number of epochs to 10, as a result of an observation that the model reached convergence on the input data within 10 epochs. We also set arguments regarding whether or not to run training on the training set, evaluation on the validation set, and predictions on the test set all to “True”. Notably, our code was executed on a system with a GPU, as detailed in Section 6.1.
6.1.4 Evaluation Metrics
We utilize Accuracy, Precision, Recall, and F1-score to evaluate our classifiers. Accuracy calculates the percentage of all notebooks that are predicted correctly (as either GCU or NCU) out of all the existing notebooks in our ground truth. Precision calculates the percentage of notebooks that are predicted correctly as GCU among all notebooks predicted as GCU (whether correct or not). Recall calculates the percentage of notebooks that are predicted correctly as GCU among all notebooks that are labeled as GCU in our ground truth. The F1-score assesses the harmonic mean of precision and recall. Since both GCU and NCU are equally important, the F1-score serves as an appropriate evaluation metric that demands our consideration.
Another commonly employed metric in machine learning for evaluating binary classification is the AUC-ROC curve. It summarizes the classifier’s ability to distinguish between the positive and negative classes across all possible threshold values. A higher AUC value indicates better discrimination power of the model. An AUC of 0.5 indicates random guessing, while an AUC of 1 indicates perfect classification.
In order to assess inter-rater agreement in our human labeling task among three raters, we use Fleiss’ kappa Fleiss1971Measuring . Fleiss’ kappa is a statistical measure used to assess the agreement among multiple raters when categorizing items into two or more categories. It is an extension of Cohen’s kappa, which is designed for two raters.
In research, Fleiss’ kappa is commonly interpreted similarly to Cohen’s kappa. It ranges from 0 to 1, where:
  • Values less than 0 indicate agreement worse than chance (very poor agreement).
  • 0 indicates agreement that is no better than chance (poor agreement).
  • Values between 0.01 and 0.20 indicate slight agreement.
  • Values between 0.21 and 0.40 indicate fair agreement.
  • Values between 0.41 and 0.60 indicate moderate agreement.
  • Values between 0.61 and 0.80 indicate substantial agreement.
  • Values between 0.81 and 1 indicate almost perfect agreement.
  • 1 indicates perfect agreement among raters.
6.1.5 Feature Importance Method in Machine Learning
Feature importance is a technique employed in data science to identify the most influential features or variables affecting a specific outcome. This information proves valuable when constructing predictive models or determining which features warrant deeper analysis. Various methods exist for calculating feature importance, such as permutation importance, mean decrease impurity, and SHAP (SHapley Additive exPlanations) values. Among the three methods, SHAP values are model agnostic, provide Interpretability for both global and local features, account for Feature Interactions, and are particularly useful for explaining the outcomes of machine learning models. In essence, they offer an estimation of each feature’s contribution to the predicted outcome, considering interactions between features LundbergL2017Unified ; Scott2018Consistent .
In the context of SHAP Beeswarm boxplots, each sample is represented by a point within each row of the plot. The horizontal placement of the point on the x-axis corresponds to the associated SHAP value. The density of points along each feature row reflects the strength of the relationship between the feature and the model’s output. The range of colors, transitioning from blue to red, indicates the progression of metric values from minimum to maximum.
6.2 Experiment Results
After setting up the evaluation, in this section we will present the results after the experiment. After each experiment, we will describe the quantitative and qualitative results.
6.2.1 Evaluation of the CU Criteria
We first present the results of five different methods for measuring CU:
  1. 1.
    The upvote approach: In this approach, the quantity of upvotes bestowed upon a notebook is regarded as a metric for assessing CU. This method is detailed in Section 6.1.3.
  2. 2.
    The CodeBERT model fine-tuned on upvotes: In this method, we utilize the model explained in Section 6.1.3.
  3. 3.
    The PLBART model fine-tuned on upvotes: Also in this approach, we utilize the model explained in Section 6.1.3.
  4. 4.
    UOCU approach: As described in Section 5.2, this method uses users’ opinions as a criterion for CU.
  5. 5.
    The hybrid approach: In this approach, we remove outlier data, scale the scores of approach 1 and 4, and then consider their sum as a criterion.
Table 4: CU Score Evaluation
Score F1-score Accuracy
1 Upvote Approach 0.55 0.59
2 CodeBERT fine-tuned by Upvote 0.61 0.62
3 PLBART fine-tuned by Upvote 0.65 0.66
4 UOCU Approach 0.66 0.70
5 Hybrid Approach 0.86 0.87
The outcomes generated from the five aforementioned methods are evaluated against the ground truth dataset (see Section 6.1.2), with detailed findings summarized in Table 4. As shown, the hybrid method achieves the highest Accuracy and F1-score. While the UOCU approach on its own demonstrates superior values in terms of F1-score and Accuracy compared to the previous methods, its amalgamation with users’ upvotes yields even better results. Consequently, we incorporate the hybrid method as a label within our machine learning operations.
{mdframed}
RQ1. How accurately is it possible to provide a criterion for measuring code understandability of computational notebooks based on their metadata?
Answer: Based on the results we achieved, it is possible to predict the CU scores of the notebooks on the Kaggle to a great extent based on their metadata. With a hybrid approach based on the UOCU score and upvoted notebooks, we achieved an F1-score of 86% and accuracy of 87% in predicting CU category, which represents a significant improvement over other baselines.
6.2.2 Evaluating the Predictive Model
Having evaluated our CU scores in the preceding section, we can now employ these scores as labels within our extensive dataset to facilitate the training and evaluation of our models.
Following the procedures outlined in Section 5.2 for calculating notebook CU scores in binary classification mode, in order to enhance the accuracy of classifying notebooks into GCU and NCU categories, we omitted the second and third quartiles (the middle two quartiles). As a result, a total of 66,360 (out of 132,723) notebooks remained in the dataset, comprising 33,180 notebooks labeled as NCU and 33,180 notebooks labeled as GCU. In preparation for training, we divided these notebooks into three distinct sets: the training set, the evaluation set, and the test set, with ratios of 70%, 20%, and 10% respectively. In the training phase, we employed the evaluation set to mitigate overfitting and utilized the test set to optimize the hyperparameters of our candidate machine learning models. Ultimately, to determine the optimal model, we assessed the trained models using our dataset of 1,050 ground truth notebooks as introduced in Section 6.1.2. All numerical results presented for machine learning algorithm outcomes in binary classification are based on this ground truth dataset.
Before delving into the examination of the accuracy levels in predicting binary classes using machine learning algorithms, it’s worth exploring the accuracy levels of these algorithms across two different modes: four-level classification, and three-level classification.
In the quaternary classification, we reintroduced the two sets of notebooks that were excluded during binary classification as two additional classes. The prediction results for this classification are outlined in Table 5.
Table 5: Machine Learning Algorithm Results for Quaternary Classification
Algorithm Accuracy Precision Recall F1-score
CatBoost 0.39 0.38 0.40 0.39
XGBoost 0.37 0.36 0.37 0.37
DecisionTree 0.30 0.30 0.30 0.30
Random Forest 0.38 0.36 0.39 0.36
In the context of ternary classification, we evenly divided the data into three equal parts based on their CU scores. You can find the evaluation metric results in Table 6.
Table 6: Machine Learning Algorithm Results for Ternary Classification
Algorithm Accuracy Precision Recall F1-score
CatBoost 0.49 0.49 0.50 0.49
XGBoost 0.48 0.49 0.49 0.48
DecisionTree 0.40 0.41 0.41 0.41
Random Forest 0.49 0.48 0.50 0.49
Now that we’ve assessed the prediction accuracy results for the three scenarios mentioned earlier, we can shift our focus back to evaluating the prediction accuracy specifically for binary classification.
The prediction results of the evaluation for our final dataset using four machine learning algorithms are presented in Table 7.
Table 7: Machine Learning Results for Binary Classification (on hybrid approach)
Algorithm Accuracy Precision Recall F1-score AUC-ROC
CatBoost 0.79 0.77 0.77 0.77 0.84
XGBoost 0.84 0.83 0.83 0.83 0.89
DecisionTree 0.82 0.83 0.84 0.83 0.82
Random Forest 0.89 0.88 0.89 0.88 0.94
As shown in Table 7, Random Forest produces the best results with an F1-score of 88% and an Accuracy score of 89%. We also utilize the AUC-ROC criterion, as introduced in Section Section 6.1.4, to provide additional validation for our experiments. Notably, this criterion yields a value of 94% for the Random Forest algorithm, reinforcing our confidence in its selection.
As far as we know, no prior studies have been conducted on measuring notebook CU using machine learning methods. Hence, we established a baseline relying on notebook upvotes and trained four machine learning models accordingly. The outcomes are presented in Table 8. Notably, training based on the UOCU criterion has shown substantial advancements in accuracy and F1-score.
Table 8: Baseline ML Results for Binary Classification (on upvotes)
Algorithm Accuracy Precision Recall F1-score
CatBoost 0.76 0.71 0.74 0.74
XGBoost 0.84 0.74 0.78 0.80
DecisionTree 0.76 0.72 0.74 0.75
Random Forest 0.82 0.75 0.79 0.80
{mdframed}
RQ2. Is it feasible to develop a model capable of achieving satisfactory performance in assessing the code understandability of notebooks?
Answer: We developed a predictive model using the Random Forest machine learning algorithm to evaluate CU in notebooks through binary classification. This endeavor yielded a commendable accuracy of 89% along with an F1-score of 88%, showcasing significant improvements in accuracy and F1-score compared to the baseline.
6.2.3 Finding Features Effective on CU
In this study, we aimed to investigate the factors that impact CU in a dataset of notebooks. To achieve this, we applied the SHAP framework to assess the importance of different features, as discussed in Section 6.1.5. Our results are presented in Figure 4, which illustrates a Beeswarm summary box plot. This plot offers a condensed overview of how the most significant notebook features impact the model’s output.
Figure 4: The Beeswarm Boxplot for SHAP Values
Figure 4 of SHAP values reveal that performance tier (PT) is the most important feature, on average, in determining CU. Users with lower performance tiers are generally less likely to produce understandable notebooks. Furthermore, the higher the expertise, the cleaner the notebook code. This result was consistent across our experiments with different learning algorithms and various measures of CU. The takeaway is straightforward: investing in training and enhancing the notebook developer experience is crucial for achieving a better understanding of the code.
Perhaps we are looking for a model capable of predicting the CU score using static metrics extracted from a notebook, regardless of the developer’s skill level. In this context, assigning a value to the PT metric might not be straightforward. In a separate experiment, we omitted this metric and attained an accuracy and F1-score of 80% using the Random Forest algorithm.
According to the SHAP chart, the next most important features in determining CU are a script-based metric and four notebook-based metrics as follows:
  • Number of blank lines of code (BLC), which is related to notebook code, has a negative impact on CU.
  • Visualization data types (VDT), which is related to code cell output, has a positive impact on CU.
  • Number of executed cells (EC), which is related to code cell output, has a positive impact on CU.
  • Number of H1 tags in markdown headlines (H1), which is related to markdown cells, has a positive impact on CU.
  • Number of lines of markdown cells (LMC), which is related to markdown cells, has a positive impact on CU.
Table 9: The Importance of Each Group of Notebook Metrics in CU (on hybrid approach)
metric group Best Algorithm Precision Recall F1-score Accuracy
notebook-based Random Forest 0.85 0.83 0.84 0.84
script-based Random Forest 0.77 0.76 0.76 0.76
basic metrics Decision Tree 0.77 0.76 0.76 0.76
complexity metrics Decision Tree 0.75 0.74 0.74 0.75
readability metrics Decision Tree 0.75 0.76 0.75 0.76
Halstead metrics Decision Tree 0.77 0.76 0.76 0.76
We continued to investigate the impact of each group or subgroup of code metrics on code comprehensibility. The results of this investigation are provided in Table 9. As evident, notebook-based metrics are more effective on their own compared to script-based metrics. Moreover, within the category of script-based metrics, basic and halstead metrics demonstrate greater effectiveness.
{mdframed}
RQ3. Which notebook metrics can affect CU, and to what extent can each metric influence code understandability?
Answer:After training our models based on CU scores, we used Shapley scores to uncover the most impactful features of CU. At a broader level of analysis, notebook-based metrics exhibit the strongest influence on CU. Drilling down further, metrics like the number of blank lines of code, number of visualization data types, number of executed cells, number of H1 tags in markdown headlines, and number of lines of markdown cells all play a crucial role.
7 Related Work
This section provides an overview of related work in the areas of “Jupyter Notebooks Code Quality” and “Measuring Code Understandability Based on Opinion”.
7.1 Jupyter Notebooks Code Quality
High-quality code can expedite the process of code comprehension. Several studies related to code quality in Jupyter notebooks have been done in recent years. In a qualitative study, Dong et al. dong2021splitting investigated how scientists clean their code on Jupyter notebooks. They sampled 20,000 notebooks with more than ten commits and classified them into three categories: personal, sharing, and productive notebooks. Finally, they sampled ten notebooks from each group and analyzed the changes between subsequent commits. Based on their findings, deleting commented blocks, adding commented blocks, reordering cells, adding markdown, renaming variables, splitting code into multiple Python files, and reorganizing code into functions and classes are among the most frequent activities done by data scientists to clean their code.
One of the activities that programmers engage in during the code comprehension phase of programming is executing different parts of the code and monitoring state variables. In a study by Wang et al. wang2020assessing , 10,000 GitHub repositories containing Jupyter notebooks were randomly cloned. They found that 73% of the repositories were not reproducible, which can have a negative impact on code understanding.
Venkatesh et al. introduced an approach aimed at enhancing comprehension and navigation within Jupyter notebooks venkatesh2023enhancing . The paper addresses the challenges associated with undocumented notebooks. The authors propose a tool named HeaderGen to tackle these challenges. HeaderGen automatically supplements code cells with annotations derived from a taxonomy of machine learning operations. This augmentation aims to enhance both comprehension and navigation within notebooks. In a user study, participants reported finding HeaderGen beneficial.
In a study by Zhu et al. zhu2021restoring , over 4,000 Kaggle notebooks were sampled to investigate the causes of a large portion of repositories being unexecutable. The study found that the use of deprecated APIs was the primary cause of 31% of unexecutabilities. To address this issue, they developed a library called Relancer that can automatically diagnose, infer, and upgrade deprecated APIs using a divide and conquer approach.
While the ability to execute code cells in different orders can help data scientists try out different settings and find the best one, it also presents a challenge for the reproducibility of results. The order of code cells in a notebook may not match the order in which the original developer executed them. To address this issue, Wang et al. developed a tool that can automatically find the possible orders that satisfy the dependencies between code cells wang2020assessing .
Wang et al. wang2020better used the PEP8 checker to analyze the code quality in Jupyter notebooks shared in two Kaggle competitions. The PEP8 checker validates whether Python code follows Python coding style best practices, such as line length and spacing between operators. They observed that 36.26% of lines of code in Jupyter notebooks led to PEP8 checker errors, which may negatively affect code understanding. In contrast, when the PEP checker was applied to another dataset of Python scripts, only 13% of Python lines were erroneous. This study clearly demonstrates that the quality of Jupyter notebook code is significantly lower than that of regular Python scripts based on PEP checker metrics. Grotov et al. grotov2022large conducted a comparison between Python code written in Jupyter notebooks and traditional Python scripts, examining them from two perspectives: structural and stylistic. In terms of structure, they discovered several distinct aspects in which notebook code differs from its traditional counterpart. Notebooks exhibited a unique pattern in the utilization of functions, and the complexity metrics indicated that they contain code that is structurally simpler but more intertwined. This could potentially decrease their comprehensibility. In terms of stylistic differences, the disparities were more pronounced, with notebooks having 1.4 times more stylistic errors compared to scripts.
Literate programming is a programming paradigm that involves adding human-understandable natural language to small code snippets to make them as easy to understand as possible knuth1984literate . Pimentel et al. pimentel2019large used the GitHub API to analyze GitHub repositories between 2013 and 2018, and found that almost all Jupyter notebooks had markup cells, which is a characteristic of literate programming.
Wang et al. wang2022documentation compared a limited number of notebooks (100 notebooks) with high upvotes to a large dataset of randomly selected notebooks on Github. After observing that the notebooks with higher scores had an average number of markdown lines greater than the other group, they concluded that markdown cells and documentation play a significant role in the code quality of Jupyter notebooks. The authors then implemented a model named Themisto, which can automatically generate documentation for code cells. Like this study, our study also considers the number of documentation lines as one of the possible metrics that can affect code quality. However, unlike Wang et al.’s study, which was based on a limited number of notebooks, we evaluate this metric across a large number of notebooks in our dataset.
The studies mentioned attempted to measure code quality using various features derived from Jupyter notebooks, but none of them measured the extent to which each feature can affect CU. Some of these studies only analyzed a small number of notebooks, which may not be representative of the broader population of Jupyter notebooks on GitHub. There is currently no widely accepted set of metrics for evaluating Jupyter notebook code quality. Different studies may use different metrics, making it difficult to compare results across studies. Our study is different and complements these studies by applying a machine learning model to determine if each metric is correlated with CU. We also report which features have a significant impact on code comprehension and which features have an insignificant effect.
7.2 Measuring Code Understandability Based on Opinion
In this section, we provide background information on studies related to our research, which aims to develop a comprehensive metric for understandability based on software developers’ opinions.
Olivera et al. oliveira2020evaluating categorized the criteria used in a code comprehension study into five groups: correctness, time, opinion, visual metrics, and brain metrics. According to their findings, over half of the studies used people’s opinions as a measure of CU. These studies relied on the subjective preferences and instincts of the participants instead of analyzing their actions or performance outcomes. In the following section, we will review some of these studies.
One such study is by Sykes et al. sykes1983effect , which investigated the level of programmers’ understandability of the start-end states of blocks in Pascal programs. To achieve this, the study developed two types of questionnaires - a subjective questionnaire that asked people to express their understanding of the program in one sentence, and a code comprehension test with ten questions from different parts of the Pascal code.
Buse and Weimer buse2009learning explored the concept of code readability and investigated its relation to software quality. They collected data from 120 human annotators to derive associations between a simple set of local code features and human notions of readability. The annotators rated Java code snippets on a scale of one to five, where a higher score represented greater readability.
Medeiros et al. medeiros2018investigating aimed to investigate the relevance of misunderstanding patterns in C code by analyzing 50 open-source C projects on GitHub. They used a combination of repository mining and a survey to identify common misunderstanding patterns and evaluate their impact on understanding the source code. The survey presented several code patterns to participants and asked them about the level of influence each pattern had on their understandability of the code.
Scalabrino et al. scalabrino2019automatically conducted an extensive evaluation of 121 existing, as well as new, code-related, documentation-related, and developer-related metrics. They attempted to correlate each metric with understandability and build models that combine metrics to assess understandability. To do this, they used 444 human evaluations from 63 developers. One of the criteria they used to measure CU is called Perceived Binary Understandability (PBU), which is based on developers’ opinions. This is a binary categorical variable that is true if a developer perceives that they understood a given code, and false otherwise. Finally, they obtained a bold negative result: none of the 121 experimented metrics is able to capture CU. As in the previous article, the methodology of this study is based on collecting opinions from a limited number of developers. The research scope of this article was limited to Java-based software. PBU criteria were also used in scalabrino2019automatically for CU measurement.
The challenges of these studies are that they relied on collecting opinions from a limited number of developers or on human labeling for the training dataset, respectively. Collecting opinions from a limited number of developers could limit the generalizability of the findings, as the opinions may not be representative of a broader population of developers. On the other hand, relying entirely on human labeling for the training dataset could limit the scalability of the approach, as it is not feasible to manually label large amounts of code.
There are some studies on software repositories that assess the quality and understandability of the code using factors like user ratings or votes. For instance, Nasehi et al. nasehi2012makes analyzed a large dataset of code snippets from StackOverflow and identified several features of high-quality code examples based on various evaluation metrics, including the number of upvotes, the number of views, code complexity, and the presence of error handling. Based on their analysis, they concluded that user feedback, as reflected by upvotes and downvotes, is a significant factor in determining the quality of code examples on StackOverflow.
Lu et al. lu2018internal consider user stars as a measure for identifying high-quality code in GitHub, although it was not the main focus of the study. The authors mentioned that stars are an indication of a project’s popularity, and that a high number of stars can attract more developers to contribute to a project.
According to some researches wang2022documentation ; Wang2021WhatMakes ; mondal2023cell2doc , community voting is a good indicator of a computational notebook’s quality. When community members upvote a notebook on Kaggle, they believe they are voting on the readability and completeness of the computational narrative. Liu et al. liu2021haconvgnn built their dataset around this hypothesis. They categorized notebooks with more votes on Kaggle as notebook with better code quality and well documentation.
The main challenges posed by the approaches presented in these studies are that they solely rely on user votes on software repositories. These metrics may be affected by factors beyond code comprehensibility, including ease of use, algorithm accuracy, popularity, and marketing considerations.
8 Threats to Validity
This section details potential threats to the validity and limitations of our findings, alongside our approaches to mitigate them.
8.1 Internal Validity
Internal validity refers to the degree to which any external factors or unanticipated biases may have impacted the design and analysis of the study. This includes confounding variables that could have affected the results easterbrook2008selecting .
A primary concern regarding our research’s validity is the feasibility of determining the CU of notebooks based on user comments. This research hinges on our capability to automatically tag each notebook with an understandability label derived from these comments. To validate this approach, we consulted four experts, asking them to annotate the comments in relation to CU. The analysis revealed that a significant portion of comments indeed align with CU metrics. Furthermore, to determine if machine learning models could discern this relationship, we trained a model to predict the relevance of comments to CU; the model demonstrated promising results on test data.
To ensure notebooks garnered enough user engagement to gather varied opinions, we set a baseline of 500 views. Notebooks falling short of this mark were excluded from our analysis.
8.2 External Validity
External validity refers to how well the research questions align with the research objectives and the degree to which conclusions can be applied beyond the specific study easterbrook2008selecting .
While much of the existing research has evaluated CU primarily on the basis of code features, we identified that in Jupyter notebooks, the accompanying documentation plays a pivotal role in influencing CU. Recognizing this, we augmented our feature set, extending beyond the sole focus on code-related attributes, to include documentation-related aspects as well. As a result, our findings hold relevance not just for traditional coding environments such as Java or C, but notably for environments like Jupyter notebooks where markdown cells and documentation significantly bolster CU. This is further exemplified by how code comments, even within brief code snippets, can mirror the functionality of markdown cells in elevating understandability.
Even though we expanded the code criteria used by similar previous studies, we bounded our analysis to primitive criteria affecting CU. Motivated by low code quality in Jupyter notebooks and considering their important role for educational purposes, as the first study investigating CU in Jupyter notebooks, our primary goal was to take the first step toward devising a comprehensive metric for assessing CU in Jupyter notebooks. So, we bound our study to primitive code criteria and left the rest to future studies.
8.3 Construct Validity
Construct validity is defined as an assessment of whether the theoretical constructs are being accurately understood and measured. This involves testing whether the interpretation of the constructs aligns with their intended meanings easterbrook2008selecting .
To ensure construct validity, each feature in our dataset is extracted based on the procedure explained in Section 3. Other members review the implementation by each author, and the errors are communicated to the developers. Even though we used a peer-review process to minimize the risk of failures in our implementations, It is still possible for some logical errors to remain unseen.
8.4 Replicability
We presented the experimental setup through multiple Jupyter notebooks. Moreover, the dataset and source code for training and evaluation of our models are publicly available online, allowing fellow researchers to replicate our results.
9 Conclusions and Future Work
In this paper, we have introduced a novel metric for evaluating CU in computational notebooks, which leverages user comments. Our proposed methodology integrates the DistilBERT model to identify pertinent comments, incorporating upvotes and visits to determine CU. Notably, our measure has exhibited superior performance compared to four baseline models, as validated through a survey conducted by 42 experts who assessed 1,050 notebooks. Additionally, we gathered 34 metrics from a dataset of 132,723 notebooks optimized for machine learning operations, achieving an 89% accuracy and an 88% F1-score in predicting the impact of these metrics on CU via binary classification with the Random Forest classifier. SHAP charts revealed the most influential metrics to be the “developer performance tier”, “number of blank lines of code”, “number of visualization data types”, “number of executed cells”, “number of H1 tags in markdown headlines”, and “number of lines of markdown cells”, all playing a crucial role on CU. As a matter of fact, while upvotes, traditionally utilized in prior studies for assessing CU, remain valuable, our results highlight that incorporating user comments and view counts can enhance the accuracy of this assessment. Consequently, based on our criterion, notebooks can be effectively classified with heightened accuracy utilizing the introduced notebook and code metrics.
Data Availability Statement.
Our dataset and experimental source code are available online at https://github.com/ISE-Research/NotebookCU.
Moreover, our service, which accepts a notebook and returns 33 code metrics and predict its CU level, is accessible at http://46.249. 101.150/docs.
Conflict of Interest.
The authors declared that they have no conflict of interest.
Appendix A Distribution of Comments
Table 10: A Sample of Comments Labeled with Experts
Kaggle_Id Comment Positive CU CU-unrelated Negative CU
388391 Great score! 1
1296638 Great work! Could you please checkout my work too. Thanks!
https://www.kaggle.com/aditya…
1
1165527 One of the best notebook resources in Kaggle, appreciated. Upvoted. 1
654526 thanks~ 1
724117 Excellent Notebook with interactive visualizations. Sometimes algorithms and ML are so funny. That’s sarcasm. I’m 55 in Notebook and Dataset 254. Both are great. 1
1574460 You changed my view toward data cleaning and also analyzing data. Cool Notebook! 1
1430133 Thanks kaggle! I wanted to start working with kaggle for months. This was a pretty nice way to start and finally I did it :)) 1
3181480 I don’t understand the point of sharing a blend kernel, if you’re not going to explain what you’re blending. All I see is that you have a black box which produces 1.448 LB score 1
1327863 Good job @artaxae! This notebook is very informative. You have explained all the data processes in a detailed way. 1
310390 thank you 1
1734304 clean notebook! Thanks for sharing 1
1596760 Nice example to get started in the competition 1
885042 Well explained Ashok 1
1436169 Nice work! I have a similar work based on emotion recognition from tweets. I would like to have your feedback on this notebook. 1
921444 Nice Work. Notebook was very simple to follow !!! 1
605630 Thank you for sharing! I learned so many things through this kernel. 1
1023754 thank you for your work 1
577715 Thanks for the sharing, learn a lot:) 1
1320208 Thanks for sharing. I am a beginner in Python. These step by step explanations are very much useful for me. 1
825456 wow, congrat! 1
1454626 It was my first time using Random Forest! So much to learn 1
882902 Really helpful notebook to learn ML, thanks for sharing your knowledge. 1
Appendix B Performance Tier Calculation
The performance tier in Kaggle is divided into five levels, each reflecting different levels of expertise and achievement on the platform. Here are more details about each performance tier:
  • Novice: The Novice tier is the starting point for new users. Users typically start at this level when they join Kaggle and have not yet participated in competitions or completed tasks to earn points. Novice users are encouraged to explore the platform, learn, and gradually progress to higher tiers.
  • Contributor: As users actively participate in Kaggle competitions and tasks, earn points, and contribute to the community, they progress to the Contributor tier. Contributors demonstrate a commitment to improving their skills, engaging with the Kaggle community, and making meaningful contributions to various projects.
  • Expert: The Expert tier represents a significant milestone in a user’s Kaggle journey. Users in the Expert tier have achieved a high level of skill and experience in data science and machine learning. They have demonstrated consistent success in competitions, gained recognition for their work, and actively share their knowledge with others in the community.
  • Master: Users who reach the Master tier are among the top performers on Kaggle. Masters have a proven track record of success in competitions, often winning or ranking highly in challenging tasks. They are respected for their expertise, innovation, and ability to solve complex data problems effectively. Being a Master signifies a high level of proficiency and accomplishment in data science.
  • Grandmaster: The Grandmaster tier is the highest level of achievement on Kaggle. Grandmasters are elite data scientists who have reached the pinnacle of success on the platform. They are highly skilled, innovative, and influential in the data science community. Grandmasters are admired for their exceptional performance, groundbreaking work, and leadership in the field.
Each tier represents a progression in a user’s skills, knowledge, and impact within the Kaggle community. Advancing through the tiers requires dedication, hard work, continuous learning, and a passion for data science.